Suppose the demand curve for a product is P=12-.5Q. Which of the following is accurate?
12. Suppose the demand curve for a product is P 12-0.50. Which of the following is accurate? a. The maximum willingness to pay for the 10th unit is $7. b. If the price is $8, the quantity demanded is 4 units. c. The maximum willingness to pay for the 8th unit is $10. d. If the price is $6, the quantity demanded is 10 units. e. None of the above.
12. Suppose the demand curve for a product is P- 12-0.50. Which of the following is accurate? a. The maximum willingness to pay for the 10h unit is $7. b. If the price is $8, the quantity demanded is 4 units. c. The maximum willingness to pay for the 8th unit is $10. d. If the price is $6, the quantity demanded is 10 units. e. None of the above.
Answer #1

Ans12) the correct option is a) the maximum willingness to pay for the 10th unit is $7

When Quantity is 10, put Q = 10 in P = 12 - 0.5Q

P = 12 - 0.5(10) = 7

Price is the maximum willingness to pay so it is 7